OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (uri):

2x+y=19 -x+y=7 Multiply the first equation by 2 and the second equation by 4 2(2x+y=19) 4(-x+y=7) Multiply and subtract them :)